OpenStudy (anonymous):

Rewrite the equation using function notation. Graph the function and show the intercepts. My equation is y=5x+3. Im sure this is really simple i just have no clue how to do it.

OpenStudy (afifarahman):

\[\frac{ x }{ 3/5 }-\frac{ y }{ 3 }= 1\] wherre 3/5 is x intercept and 3 is y intercept.